OpenStudy (anonymous):

What is the factored form of 2x^2 + 20x + 50

OpenStudy (anonymous):

So far, I took out 2

OpenStudy (ness9630):

Okay, let's see what you got.

OpenStudy (anonymous):

And that left me with 2(x^2 + 10 x + 25)

OpenStudy (anonymous):

so its 2(x+5)(x+5)

OpenStudy (ness9630):

This can be done without the quadratic formula. And @IrisL please don't give away answers,

OpenStudy (anonymous):

I thought the answer was 2(x +5)(x + 5) too

OpenStudy (anonymous):

However, it isn't one of the listed answers

OpenStudy (anonymous):

yea exactly, its 2(x+5)(x+5)

OpenStudy (anonymous):

maybe your textbook solution is wrong, you don't need the quadratic formula when it is factorable.

OpenStudy (ness9630):

Hmm, what are your solutions listed? Just curious.

OpenStudy (anonymous):

(x + 5)(2x − 8) (x − 5)(2x + 8) (x + 5)(2x + 10) (x + 5)(2x − 10)

OpenStudy (ness9630):

Oh, simple, just distribute that 5 back in.

OpenStudy (ness9630):

Ugh, stop giving away answers!

OpenStudy (anonymous):

How would I get to that answer though? Distribute which 5?

OpenStudy (ness9630):

oops, sorry I meant 2. Okay, so we have 2(x+5)(x+5) Multiply one of them with the 2, so [2(x+5)](x+5) Do what's in the brackets.

OpenStudy (anonymous):

Oh! Okay. So it would be (2x+10)(x+5)

OpenStudy (anonymous):

Cools XD

OpenStudy (ness9630):

Yup :P

OpenStudy (anonymous):

What about polynomials of the third degree?

OpenStudy (anonymous):

64^3 +343, to be exact

OpenStudy (ness9630):

Anything you can factor out right off the bat?

OpenStudy (anonymous):

2 isn't a factor, 3 isn't a factor, 4 isn't a factor... I'm not sure.

OpenStudy (anonymous):

7 is a factor of 343, but not 64

OpenStudy (ness9630):

Just as a notice, this has one solution..

OpenStudy (anonymous):

How do you know?

OpenStudy (ness9630):

I graphed it :)

OpenStudy (anonymous):

lol cool

OpenStudy (anonymous):

So, how would I go about it without any factors? I notice they are both perfect cubes

OpenStudy (anonymous):

I think

OpenStudy (ness9630):

Oh, I feel dumb now. What we have to do is solve it as a normal equation. So set it equal to zero, and move everything away from x.

OpenStudy (anonymous):

\[64x = \sqrt{-343}\]

OpenStudy (anonymous):

But wouldn't that be bringing in imaginary numbers?

OpenStudy (anonymous):

I meant to put a three as the power

OpenStudy (ness9630):

Hold your horses there, you forgot tht 64

OpenStudy (ness9630):

*that

OpenStudy (anonymous):

lol \[x = \frac{ \sqrt[3]{343} }{ 64 }\]

OpenStudy (anonymous):

ANd I meant negative -343

OpenStudy (anonymous):

x = 7i/64

OpenStudy (anonymous):

Factor out 2, 2(x^2 + 10x + 25) = 2 (x +5)(x+5) or you can write it as 2(x+5)^2. Done.

OpenStudy (anonymous):

You can take the cube root of a negative number. It's only even powers that get into the imaginary stuff.

OpenStudy (ness9630):

Close: \[64x^3+343=0\] \[64x^3=-343\] \[x^3=-\frac{343}{64}\] \[\LARGE \sqrt[3]{-\frac{343}{64}}\] Now solve it. And when you are taking the cubic root if anything, there are no imaginary numbers, it's just negative. Because \[\LARGE -x^3=-x*-x*-x\]

OpenStudy (anonymous):

(Well, negative numbers have complex cube roots too. But I assume you care more about the real one.)

OpenStudy (anonymous):

Okay. So - 343/64 = 5.35?

OpenStudy (ness9630):

Take the cubic root of each, first.

OpenStudy (anonymous):

Oh. so 7/4

OpenStudy (ness9630):

*-7/4 but yessssssss

OpenStudy (anonymous):

lol okay. So -1.75?

OpenStudy (ness9630):

Yea, and that's your one solution.

OpenStudy (anonymous):

4x − 7 16x2− 28x + 49 16x2+ 56x + 49 None of the above

OpenStudy (anonymous):

So the first one?

OpenStudy (ness9630):

Remember what I taught you about working backwards from a solution? x=-7/4

OpenStudy (anonymous):

Yes, so far 4x + 7?

OpenStudy (anonymous):

Oh. So none of them?

OpenStudy (ness9630):

Looks like it. Because it factors into: \[\LARGE (4x+7)(64x^2-112x+196)\]

OpenStudy (ness9630):

Wait, no ignore that.

OpenStudy (anonymous):

lol k

OpenStudy (ness9630):

Hmm, you know what.. I'm stuck >.> Open a new question and see if anyone else can answer it. Sorry. If I get anything I'll let you know.

OpenStudy (ness9630):

After distributing, I got that the factors were (4x+7)(16x^2-28x+49)